# Universal Vessel Segmentation for Multi-Modality Retinal Images

### v3

@@ IV-B1 Translation Network
The balance between the size of the generator and discriminator is important to stably train the CycleGAN for our task. We use a U-Net [31] generator with 16.66M parameters and an NLP discriminator with 2.77M parameters [^2] More details on the network structures we use for translation and segmentation will be provided in our codes.. For training, we use a batch size of 4 and train the model for 100 epochs on the 4000 images. We start with an initial learning rate of 2e-4 and start decaying the learning rate linearly to 0 from epoch 50 to 100. Adam optimizer is used to train both the generators and the discriminators, with a momentum of $(0.5,0.999)$ and no weight decay. $\lambda_{c}=$10.0 and $\lambda_{i}=$0.5.
@@ IV-B2 Segmentation Network
An initial learning rate of 1e-3 is used to train the model for 150 epochs, then decayed to 1e-4 to train for another 250 epochs. Adam optimizer is used with a weight decay of 1e-3 and momentum of $(0.9,0.999)$. We train with a batch size of 1. The weights $\lambda_{tc}$ and $\lambda_{ts}$ are 0.05 and 0.0002, respectively. All the testing images are first padded to square and resized to $\times768$ to pass through the translation and the segmentation network. Then the prediction is resized back to the original image size to compute the metrics with the ground truth labels. The implementation of both image translation and the segmentation parts are under the PyTorch framework. The computation of persistent homology relies on the GUDHI library [89].
@@ IV-B3 Data Augmentation
Extensive data augmentations are used for training both the translation and the segmentation network. For translation network, random scaling, sharpness, contrast and gamma adjustment are used. Additionally, since different camera manufacturers apply different shapes of mask to the image, we randomly apply the masks shown in Fig. 7 to the images to prevent the model from generating artifacts out of the effective imaging area. For segmentation network, in addition to the aforementioned methods, we also apply random negative film followed by a contrast or sharpness adjustment. This is essential to allow the universal segmentation model to work for the FA modality as the vessels appears to be brighter than the background and we found no image translation method capable of stably inverting the color of the vessels. In addition, random image rotation, shift and Gaussian noise are applied.
IV-C Comparison with Finetuned SOTA Methods
IV-C1 UVSM on retinal images in conventional field-of-view
@@
To the best of our knowledge, this work is the first universal retinal vessel segmentation model and there is no previous method that we can fairly compare with. Therefore, we compare with state-of-the-art (SOTA) methods which are finetuned individually on each dataset either by full-supervision or unsupervised domain adaptation. Since we use three new datasets and the works in the retinal vessel segmentation literature have inconsistent dataset choice, dataset split, and inconsistent choice of evaluation metrics, we prepare a new benchmark where all the methods are evaluated in a consistent way for fair comparison. The baseline methods are run using the recommended training details in respective papers and in a five-fold experiment that splits the whole dataset by 5 groups. Each fold uses one group for testing and the other four for training. Eventually, all images are used for testing so as to make sure that the average score is computed on the same images with our universal model (which is directly tested on all the images). For the unsupervised domain adaptation methods, we use the DRIVE dataset with annotation as the source domain and the per-fold training set of each dataset as the target domain for training. The supervised method baseline includes U-Net [31], CS2-Net [33], W-Net [42], FA-Net [43], GT-DLA-dsHff [34] and WS-DMF [41]. For domain adapation, the baseline includes Pse. Lab. (pseudo-label-based)[42], FRR-TSNet (teacher-student network-based) [77], MLAL (adversarial learning-based) and MLAL-PDSF (hybrid adversarial learning and teacher-student) [76]. The training and testing of the baselines are done in original image size except for the HRF dataset. The images in HRF dataset are too large to be trained on our 24GB GPUs. Therefore, we first pad the image and ground truth to square, then downsample them to $\times768$ for training and inference. The predicted vessel likelihood is then upsampled to the original image size to compare with the original ground truth and compute the metrics.
@@
Since the use of evaluation metrics is inconsistent in the literature, in our work we adopt a consistent benchmark and use most of the commonly-used metrics, including the pixel-wise accuracy (Acc.), the Dice score, the specificity (Sp), the sensitivity (Se), the Precision (Pr), the F1 score and the Mathew Correlation Coefficients (MCC). In addition, although the main goal of topological feature learning in this work is for domain generalization, topological accuracy is still important in retinal vessel segmentation. Therefore, we also use three topological metrics which are commonly used in the segmentation of more general tubular structures. These metrics include the clDice score [65] and the 0-/1-Betti errors ($\beta_{0},\beta_{1}$) [66]. For all of the reported metrics, we compute them at the original image size and average over all test images.
@@
The results in Table II, Table III and Fig. 8 show that our proposed model, not only allowing universal segmentation, but also has comparable performance with SOTA finetuned supervised methods and outperforms the finetuned unsupervised domain adaptation methods. However, for some modalities there still exist non-neglectable performance gap after the supervised methods, especially for the FA. As shown by an example in the third row of Fig. 8, the FA modality (particularly from the Optos camera) has very high contrast between the vessel and the background. As a result, more ending vanishing vessels are visible on the image and the universal model often fails to identify them. The reasons are two-fold: (1) The image translation introduces non-avoidable detail loss; (2) The topological loss function makes the model more conservative in predicting a pixel as vessel.
@@
Figure 9: Qualitative results on Ultra-Wide-Field (UWF) retinal images. From left to right: image, translated image, UVSM segmentation, ground truth. Red arrows show the outlier patterns in the peripheral causing false segmentation, which is a current limitation of the UVSM.
@@
IV-C2 UVSM on OCTA and UWF
We separately evaluate OCTA modality and Ultra-Wide-Field retinal images (UWF) since they have considerably different field of views and the image feature in the OCTA modality is significantly different from the other modalities. In addition, vessel segmentation of these two types of retinal images has independent literature and baselines, respectively. For the metrics, we follow and use the metrics in [86] for OCTA modality and the metrics in [90] for UWF imaging.
For OCTA, we compare with SOTA supervised methods: UNet++ [91], FARGO [92], FRNet-base and FRNet [86]. We observed that the UVSM has good performance in predicting the major vessels. But the prediction for the thin ending vessels is usually missing (Fig. 8). Generally speaking, there are still notable performance gap between the UVSM and SOTA finetuned methods (Table IV). But keep in mind that we only use 19 annotated CF images to train the model, while our baseline methods are finetuned on OCTA training set with highly similar data with the testing set and are over 7 times larger. Overall, the zero-shot results on OCTA shows the strong generalization ability of our UVSM and the effectiveness of our proposed learning methods.
For UWF, we compare with SOTA supervised methods: LKM-UNet [93], Swin-Umamba [94], EM-Net [95] and Serp-Mamba [90]. Experiment shows that the UVSM has strong performance in the center area of the retina but can generate more artifacts at the boundary of the image (Fig. 9). The overall performance is comparable with the state-of-the-art methods (Table IV). The results on UWF images again shows the the strong generalization ability of our proposed model.
Figure 12: Comparison of the final segmentation performance between using different methods for image translation. Reporting metric: Dice score.
Figure 13: Qualitative examples of the translated images using different image translation methods. From top to bottom: MC, FA, FAF, IR, OCTA. (best zoom in to view)
@@ IV-D Comparison with general-purpose models
Although great progress has been made in general-purpose segmentation models in recent years, they struggle with specialized image and structures such retinal image and curvilinear structures. To further highlight the need for a universal retinal vessel segmentation model, we compare with three recent foundational vision models: SAM2 [14], Grounded SAM2 [15] and the SOTA variant for general medical image segmentation-MedSAM2 [16]. For Grounded SAM2, we use a uniform prompt word ‘vessels.’. For SAM2 and MedSAM2, we provide a manual point prompt with over 100 positives points on the vessel and over 20 negative points on the background for the first image of each dataset and propagate the mask to all following images. Due to space constraint, we report the result per modality, i.e., average the score over all images (from one or multiple datasets) from each modality. The UWF image is excluded from the MC modality since the extremely high resolution of UWF images makes it infeasible to run these models on our available hardware. The result in Table V shows that our proposed UVSM has much better performance than these general-purpose segmentation models.
@@ IV-E Ablation Studies
Due to high-volume of experimental results, for all the ablation studies, we report the result per modality the same way in section IV-D.
@@ IV-E1 Pipeline
In this part, we study how the main components in the proposed pipeline improve the segmentation quality. All the experiments in the ablation studies adopt the same training configurations, data augmentations and network structure, unless otherwise stated. We first run a baseline experiment where no image translation nor topological learning is applied. In another word, we simply train a segmentation model on the DRIVE training set and apply the model to the multi-modality retinal images. As shown by Fig. 11, the performance of the model is very poor. Both image translation and topological learning improves the performance of the model, while topological learning plays a more dominating role. Optimal performance is achieved by using both I2I and topological learning. An interesting observation is that, for the FA modality, the vessel color is inverted and the model tends to predict the boundary of the vessels as foreground rather than the vessel itself, resulting in a ‘sandwich’ pattern. None other domain adaptation method can solve this issue except for the proposed topological learning (Fig. 10).
@@ IV-E2 Image Translation Method
We compare CycleGAN with two most recent Diffusion model-based methods (StyleDiffusion [61] and UNSB [56]). The 19 translated DRIVE images are from the respective translation method to train the downstream segmentation network.
For StyleDiffusion, we randomly pick only one CF image as the style target and reduce the training set size to the same as that in [61] because we find that on larger training set, StyleDiffusion generates severe artifacts despite our efforts to finetune the training hyperparameters for our images. In addition, data augmentation cannot be implemented in StyleDiffusion for its two-diffusion-stage design. In general, the final segmentation performance by StyleDiffusion is the poorest (Fig. 12). Furthermore, similar with the observation in [12], where style transfer method is also used for retinal vessel segmentation, the transferred images tend to include many vessel-like artifacts (Fig. 13) introduced by the perceptual inductive bias, which will later result in many artifacts in the segmentation.
@@
As shown in Fig. 12, when comparing the final segmentation performance, CycleGAN demonstrates the best performance for most datasets. It is also observed that the CycleGAN translation results are visually more alike with the target CF images where the sharpness and contrast of the images are weaker than in the other modalities (Fig. 13). Furthermore, for FAF where the original images are very noisy, CycleGAN also performs better in reducing the noise and generating a smoother translated image. Additionally, UNSB fails on the OCTA modality while StyleDiffusion and CycleGAN show better generalization ability.
@@
TABLE VI: Comparison between different choices of topological loss functions. While the primary goal is to improve domain generalization, we still choose to report the topological metrics here since they are more relevant to the structural properties of the vessels.
@@ IV-E3 Topological Segmentation Loss Function
In this part we compare how learning different combinations of topological features impact the performance of the UVSM. Table VI and Fig. 14 show that, both $\mathcal{L}_{tc}$ and $\mathcal{L}_{ts}$ help with overcoming the domain gap of the non-CF modalities. It is observed that $\mathcal{L}_{tc}$ is unable to address the inverted vessel color in FA images and tend to instead predict the boundary of the vessel, which results in a ‘sandwich’ pattern. In addition, using only $\mathcal{L}_{tc}$ adds saw artifacts and discontinuities to the segmentation. $\mathcal{L}_{ts}$ is able to address the problem of FA modality and improves the discontinuities along the vessel, however, it is observed that if only $\mathcal{L}_{ts}$ is used, the vessel boundary will be too smoothed and the morphological quality of the segmentation will be poor. With a combination of $\mathcal{L}_{tc}$ and $\mathcal{L}_{ts}$, the saw and smoothing artifacts counteracts, and the challenge in predicting FA vessels is solved, producing the optimal results.
TABLE VII: Comparison of using different segmentation network structures for UVSM. Reporting metric: Dice score.
@@ IV-E4 Segmentation Network Structure
We also tested and compared multiple most recent network structures in retinal vessel segmentation as the segmentation model backbone for our UVSM, namely, FRUNet [77], MCG&BSA-Net [76], RCARU-Net [96], Wave-Net [97] and ResDoU-Net [98]. Table VII presents that with different backbones, universal retinal vessel segmentation is still achieved, which shows the effectiveness and flexibility of our proposed learning methods. In addition, with some more recent backbones, the overall performance of the UVSM is further improved.

### v4

@@ IV-B1 Translation Network
The balance between the size of the generator and discriminator is important to stably train the CycleGAN for our task. We use a U-Net [31] generator with 16.66M parameters and an NLP discriminator with 2.77M parameters [^3] More details on the network structures we use for translation and segmentation are provided in our code.. For training, we use a batch size of 4 and train the model for 100 epochs on the 4000 images. We start with an initial learning rate of 2e-4 and start decaying the learning rate linearly to 0 from epoch 50 to 100. Adam optimizer is used to train both the generators and the discriminators, with a momentum of $(0.5,0.999)$ and no weight decay. $\lambda_{c}=$10.0 and $\lambda_{i}=$0.5.
@@ IV-B2 Segmentation Network
An initial learning rate of 1e-3 is used to train the model for 150 epochs, then decayed to 1e-4 to train for another 250 epochs. Adam optimizer is used with a weight decay of 1e-3 and momentum of $(0.9,0.999)$. We train with a batch size of 1. The weights $\lambda_{tc}$ and $\lambda_{ts}$ are 0.05 and 0.0002, respectively. All the testing images are first padded to square and resized to $\times768$ to pass through the translation and the segmentation network. Then the prediction is resized back to the original image size to compute the metrics with the ground truth labels. The implementation of both image translation and the segmentation parts are under the PyTorch framework. The computation of persistent homology relies on the GUDHI library [91]. The computational cost of this work is reported in Table II.
@@ IV-B3 Data Augmentation
Extensive data augmentations are used for training both the translation and the segmentation network. For translation network, random scaling, sharpness, contrast and gamma adjustment are used. Additionally, since different camera manufacturers apply different shapes of mask to the image, we randomly apply the masks shown in Fig. 7 to the images to prevent the model from generating artifacts out of the effective imaging area. For segmentation network, in addition to the aforementioned methods, we also apply random negative film followed by a contrast or sharpness adjustment. This is essential to allow the universal segmentation model to work for the FA modality as the vessels appear to be brighter than the background and we found no image translation method capable of stably inverting the color of the vessels. In addition, random image rotation, shift and Gaussian noise are applied.
IV-C Comparison with Fine-tuned SOTA Methods
IV-C1 URVSM on retinal images in conventional field-of-view
@@
To the best of our knowledge, this work is the first universal retinal vessel segmentation model and there is no previous method that we can fairly compare with. Therefore, we compare with state-of-the-art (SOTA) methods which are fine-tuned individually on each dataset either by full-supervision or unsupervised domain adaptation. Since we use three new datasets and the works in the retinal vessel segmentation literature have inconsistent dataset choice, dataset split, and inconsistent choice of evaluation metrics, we prepare a new benchmark where all the methods are evaluated in a consistent way for fair comparison. The baseline methods are run using the recommended training details in respective papers and in a five-fold experiment that splits the whole dataset by 5 groups. Each fold uses one group for testing and the other four for training. Eventually, all images are used for testing so as to make sure that the average score is computed on the same images with our universal model (which is directly tested on all the images). For the unsupervised domain adaptation methods, we use the DRIVE dataset with annotation as the source domain and the per-fold training set of each dataset as the target domain for training. The supervised method baseline includes U-Net [31], CS2-Net [33], W-Net [42], FA-Net [43], GT-DLA-dsHff [34] and WS-DMF [41]. For domain adaptation, the baselines include Pse. Lab. (pseudo-label-based) [42], FRR-TSNet (teacher-student network-based) [77], MLAL (adversarial learning-based) and MLAL-PDSF (hybrid adversarial learning and teacher-student) [76]. The training and testing of the baselines are done in original image size except for the HRF dataset. The images in HRF dataset are too large to be trained on our 24GB GPUs. Therefore, we first pad the image and ground truth to square, then downsample them to $\times768$ for training and inference. The predicted vessel likelihood is then upsampled to the original image size to compare with the original ground truth and compute the metrics.
@@
Since the use of evaluation metrics is inconsistent in the literature, in our work we adopt a consistent benchmark and use most of the commonly used metrics, including the pixelwise accuracy (Acc.), the Dice score, the specificity (Sp), the sensitivity (Se), the Precision (Pr), the F1 score, and the Matthews Correlation Coefficients (MCC). In addition, although the main goal of topological feature learning in this work is for domain generalization, topological accuracy remains crucial for retinal vessel segmentation. Therefore, we also use three topological metrics which are commonly used in the segmentation of more general tubular structures. These metrics include the clDice score [65] and the 0-/1-Betti errors ($\beta_{0},\beta_{1}$) [66]. For all of the reported metrics, we compute them at the original image size and average over all test images.
@@
The results in Table III, Table IV and Fig. 8 show that our proposed model, not only allows universal segmentation, but also has comparable performance with SOTA fine-tuned supervised methods and outperforms the fine-tuned unsupervised domain adaptation methods. However, for some modalities there still exists non-negligible performance gap compared with the supervised methods, especially for the FA. As shown by an example in the third row of Fig. 8, the FA modality (particularly from the Optos camera) has very high contrast between the vessel and the background. As a result, more ending vanishing vessels are visible on the image and the universal model often fails to identify them. The reasons are two-fold: (1) The image translation introduces non-avoidable detail loss; (2) The topological loss function makes the model more conservative in predicting a pixel as vessel.
@@
Figure 9: Qualitative results on Ultra-Wide-Field (UWF) retinal images. From left to right: image, translated image, URVSM segmentation, ground truth. Red arrows show the outlier patterns in the peripheral causing false segmentation, which is a current limitation of the URVSM.
@@
IV-C2 URVSM on OCTA and UWF
We separately evaluate OCTA modality and Ultra-Wide-Field retinal images (UWF) since they have considerably different fields of view and the image features in the OCTA modality are significantly different from the other modalities. In addition, vessel segmentation of these two types of retinal images has its independent literature and established baselines. For the metrics, we follow and use the metrics in [86] for OCTA modality and the metrics in [92] for UWF imaging.
For OCTA, we compare with SOTA supervised methods: UNet++ [93], FARGO [94], FRNet-base and FRNet [86]. We observed that the URVSM has good performance in predicting the major vessels, whereas the prediction for the thin ending vessels is usually missing (Fig. 8). In general, there are still notable performance gaps between the URVSM and SOTA fine-tuned methods (Table V). It should be noted that we only use 19 annotated CF images to train the model, while our baseline methods are fine-tuned on OCTA training set with highly similar data with the testing set and are over 7 times larger. Overall, the zero-shot results on OCTA show the strong generalization ability of our URVSM and the effectiveness of our proposed learning methods.
For UWF, we compare with SOTA supervised methods: LKM-UNet [95], Swin-Umamba [96], EM-Net [97] and Serp-Mamba [92]. Experiments show that the URVSM has strong performance in the center area of the retina but can generate more artifacts at the boundary of the image (Fig. 9). The overall performance is comparable with the state-of-the-art methods (Table V). The results on UWF images again shows the strong generalization ability of our proposed model.
Figure 12: Comparison of the final segmentation performance between using different methods for image translation. Metric reported: Dice score.
Figure 13: Qualitative examples of the translated images using different image translation methods. From top to bottom: MC, FA, FAF, IR, OCTA. (best viewed by zooming in)
@@ IV-D Comparison with general-purpose models
Although great progress has been made in general-purpose segmentation models in recent years, they struggle with specialized images and structures such as retinal image and curvilinear structures. To further highlight the need for a universal retinal vessel segmentation model, we compare with three recent foundational vision models: SAM2 [14], Grounded SAM2 [15], and the SOTA variant for general medical image segmentation: MedSAM2 [16]. For Grounded SAM2, we use a uniform text prompt, ‘vessels.’. For SAM2 and MedSAM2, we provide manual point prompts with over 100 positive points on the vessels and over 20 negative points on the background for the first image of each dataset. The generated mask is then propagated to all subsequent images. Due to space constraints, we report the results per modality, i.e., average the score over all images (from one or multiple datasets) from each modality. The UWF image is excluded from the MC modality since the extremely high resolution of UWF images makes it infeasible to run these models on our available hardware. The results in Table VI show that our proposed URVSM achieves substantially better performance than these general-purpose segmentation models.
@@ IV-E Ablation Studies
Due to the high volume of experimental results, for all the ablation studies, we report the results per modality the same way as described in section IV-D.
@@ IV-E1 Pipeline
In this part, we study how the main components in the proposed pipeline improve the segmentation quality. All the experiments in the ablation studies adopt the same training configurations, data augmentations and network structure, unless otherwise stated. We first run a baseline experiment where neither image translation nor topological learning is applied. In other words, we simply train a segmentation model on the DRIVE training set and apply the model to the multi-modality retinal images. As shown by Fig. 11, the performance of the model is very poor. Both image translation and topological learning improve the performance of the model, while topological learning plays a more dominant role. Optimal performance is achieved by using both I2I and topological learning. An interesting observation is that, for the FA modality, the vessel color is inverted and the model tends to predict the boundary of the vessels as foreground rather than the vessel itself, resulting in a ‘sandwich’ pattern. No other domain adaptation method can resolve this issue except for the proposed topological learning (Fig. 10).
@@ IV-E2 Image Translation Method
We compare CycleGAN with the two most recent diffusion model-based methods (StyleDiffusion [61] and UNSB [56]). The 19 translated DRIVE images are produced by respective translation method to train the downstream segmentation network.
For StyleDiffusion, we randomly pick only one CF image as the style target and reduce the training set size to the same as that in [61] because we find that on a larger training set, StyleDiffusion generates severe artifacts despite our efforts to fine-tune the training hyperparameters for our images. In addition, data augmentation cannot be implemented in StyleDiffusion because of its two-diffusion-stage design. In general, the final segmentation performance by StyleDiffusion is the poorest (Fig. 12). Furthermore, similar to the observation in [12], where style transfer methods are also used for retinal vessel segmentation, the transferred images tend to include vessel-like artifacts (Fig. 13) introduced by the perceptual inductive bias, which in the subsequent step will also result in many artifacts in the segmentation.
@@
As shown in Fig. 12, when comparing the final segmentation performance, CycleGAN demonstrates the best performance for most datasets. It is also observed that the CycleGAN translation results are visually more similar to the target CF images where the sharpness and contrast of the images are weaker than in the other modalities (Fig. 13). Furthermore, for FAF where the original images are very noisy, CycleGAN also performs better in reducing the noise and generating a smoother translated image. Additionally, UNSB fails on the OCTA modality while StyleDiffusion and CycleGAN show better generalization ability.
@@
TABLE VII: Comparison between different choices of topological loss functions. While the primary goal is to improve domain generalization, we still choose to report the topological metrics here since they are more relevant to the structural properties of the vessels.
@@ IV-E3 Topological Segmentation Loss Function
In this part, we compare how learning different combinations of topological features impacts the performance of the URVSM. Table VII and Fig. 14 show that both $\mathcal{L}_{tc}$ and $\mathcal{L}_{ts}$ help overcome the domain gap of non-CF modalities. It is observed that $\mathcal{L}_{tc}$ is unable to address the inverted vessel color in FA images and instead tends to predict the boundary of the vessel, which results in a ‘sandwich’ pattern. In addition, using only $\mathcal{L}_{tc}$ adds saw artifacts and discontinuities to the segmentation. $\mathcal{L}_{ts}$ is able to address the problem of FA modality and reduces the discontinuities along the vessel, however, it is observed that if only $\mathcal{L}_{ts}$ is used, the vessel boundary will be too smoothed and the morphological quality of the segmentation will be poor. With a combination of $\mathcal{L}_{tc}$ and $\mathcal{L}_{ts}$, the saw and smoothing artifacts counteract, and the challenge in predicting FA vessels is effectively mitigated, producing optimal results.
TABLE VIII: Comparison of using different segmentation network structures for URVSM. Metric reported: Dice score.
@@ IV-E4 Segmentation Network Structure
We also tested and compared several recent network structures in retinal vessel segmentation as the segmentation model backbone for our URVSM, namely, FRUNet [77], MCG&BSA-Net [76], RCARU-Net [98], Wave-Net [99] and ResDoU-Net [100]. Table VIII shows that with different backbones, universal retinal vessel segmentation is still achieved, which shows the effectiveness and flexibility of our proposed learning methods. In addition, with some more recent backbones, the overall performance of the URVSM is further improved.
IV-F Pilot Study on Topological Domain Adaptation
To further demonstrate the potential of topological learning as a domain adaptation strategy for more general curvilinear structure segmentation, we conducted a preliminary experiment beyond retinal vessel segmentation. Specifically, we directly applied the retinal vessel segmentation model trained with the proposed topological learning (corresponding to “topo” in Fig. 11) and compared it with a baseline model trained without topological learning (“vanilla” in Fig. 11). As reported in Table IX and Fig. 15, the segmentation model trained with the joint topological learning demonstrates stronger generalization capability than its baseline counterpart.
Figure 15: Qualitative examples for the pilot study.
TABLE IX: Performance of retinal vessel segmentation models with/w.o. topological learning on more general curvilinear structures
Dataset Roads CREMI
[…]
w.o. Topo 91.02 12.81 12.05 86.68 7.55 7.41
with Topo 92.44 24.20 23.19 87.06 12.02 11.89